# ———————————————————————————————————–

### v2

@@ VII Pattern counting estimator
    set the absorber parameter by using the displaced-null measurement technique developed in Girotti et al. (2024) and run the system-absorber dynamics for the remainder of the time;
@@
The first step of the adaptive protocol is to use $\tilde{n}:=n^{1-\epsilon}\ll n$ output units to produce a rough preliminary estimator $\tilde{\theta}_{n}$. This can be done by performing a repeated standard basis measurement on the output (without using an absorber). Typically, the estimator $\tilde{\theta}_{n}$ will have variance scaling with the standard rate $\tilde{n}^{-1}=n^{-1+2\epsilon}$, and $\tilde{n}^{1/2}(\tilde{\theta}_{n}-\theta)$ will satisfy the central limit theorem and a concentration bound ensuring that $|\tilde{\theta}_{n}-\theta|=O(n^{-1/2+\epsilon})$ with high probability. For instance, one can define $\tilde{\theta}_{n}$ by equating the empirical counting rate with the theoretical rate $n_{\tau}:={\rm Tr}(\rho^{\rm ss}_{\tau}K_{\tau,1}^{*}K_{\tau,1})$, see Catana et al. (2015); Girotti et al. (2023).
In the second step we set the absorber at a parameter value $\theta_{0}$ and run the system-absorber quantum Markov chain for the reminder of the time $n^{\prime}=n-\tilde{n}$. The naive choice for $\theta_{0}$ is our best guess $\tilde{\theta}_{n}$ about $\theta$, based on the first stage measurement. However, with this choice, the counting measurement suffers from the non-identifiability issue described in section II. This can be resolved by further displacing the absorber parameter by an amount $\delta_{n}:=\tau_{n}/\sqrt{n}$ where $\tau_{n}=n^{3\epsilon}$ so that $\theta_{\rm abs}:=\tilde{\theta}_{n}-\delta_{n}$. As usual we write $\theta=\tilde{\theta}_{n}+u_{n}/\sqrt{n}$ where $u_{n}$ is a local parameter satisfying $|u_{n}|\leq n^{\epsilon}$, so that $\theta=\theta_{\rm abs}+(u_{n}+\tau_{n})/\sqrt{n}$.
@@
We now describe the construction of the estimator from the outcomes of this last stage measurement. The estimator will be built using the “pattern extraction” function $\{N_{\alpha,n}(\omega):\alpha\in\mathcal{P}\}\in\mathbb{N}^{\mathcal{P}}$ that we defined in the previous section; we recall that it associates to each trajectory $\omega\in\{0,1\}^{n}$ a set of pattern counts, which is uniquely determined by the condition that $\omega$ is a maximal union of contiguous patterns separated by sequences of 00s of length at least $n^{\gamma}$ with a fixed $10<\gamma<$ 1, cf. Figure 1 d) for illustration. This means that the algorithm will not detect any pattern which contains a sequence of zeros of length larger than $n^{\gamma}$, since this would be seen as being made up of several identified patterns. We now introduce the final estimator using an intuitive argument based on extrapolating the results of Theorem 3 from fixed to slowly growing local parameters $u_{n}+\tau_{n}$. This means that $N_{\alpha,n}(\omega)$ is approximately distributed as ${\rm Poisson}((u_{n}+\tau_{n})^{2}|\mu_{\alpha}|^{2})$, for large $n$; since $\tau_{n}=n^{3\epsilon}$ is larger than $u_{n}=O(n^{\epsilon})$, the intensity of the Poisson distribution diverges with $n$, and the distribution can be approximated further by the normal $N((u_{n}+\tau_{n})^{2}|\mu_{\alpha}|^{2},(u_{n}+\tau_{n})^{2}|\mu_{\alpha}|^{2})$ with the same mean and variance. Using
@@
At the moment we only have a rigorous proof of this statement assuming a stronger version of Proposition 2, which we were not able to obtain; however, we point out that Theorem 2 in Girotti et al. (2024) establishes a similar optimality result in the case of multi-parameter estimation with independent, identical copies.
@@
The proof of the following theorem and the relationship between the extra hypothesis in Eq. (25) and Proposition 2 can be found in Appendix G.
@@
VIII Numerical experiments
In this section we illustrate the estimation protocol through numerical simulations, using a qubit model inspired by the previous work Girotti et al. (2024). The simulations are implemented in Python using the QuTiP package Johansson et al. (2013).
@@ VIII Numerical experiments
In a first simulation study we verify the predictions made in Corollary 3. We run the dynamics for a total of $n=6\times 10^{5}$ time steps, with absorber parameter $\theta_{\rm abs}=\theta_{0}$ (cf. section III.1) and system parameter $\theta=\theta_{0}+u/\sqrt{n}$ such that $\theta=0.2$ and the local parameter is $u=2$.
For each run we perform repeated measurements in the standard basis of the output units, to produce a measurement record $\omega=(\omega_{1},\dots,\omega_{n})$. For each such trajectory, the pattern counts $N_{\alpha,n}(\omega)$ are obtained by identifying patterns (sequences starting and ending with a 1) which are separated by at least $n^{\gamma}$ 00s and no pattern contains more than $n^{\gamma}$ 00s, where $\gamma>0$ is a small parameter. This can be done by combing through the sequence and identifying occurrences of a given such pattern padded by $n^{\gamma}$ 00s to the left and right (taking care of the special case of the first and last patterns). Note that for any given $n$ this procedure will not count patterns with more than $n^{\gamma}$ successive 00s. Since the mean counts for each pattern $\alpha$ is $|\mu_{\alpha}|^{2}u^{2}$ and $|\mu_{\alpha}|^{2}$ decays exponentially with $|\alpha|$, we find that patterns of such length are unlikely to occur for large $n$. The results of $N=2000$ independent repetitions of the experiment are illustrated in Figure 2 which shows a good match between the counts histograms corresponding to several patterns (in blue) and the theoretical Poisson distributions ${\rm Poisson}(|\mu_{\alpha}|^{2}u^{2})$ (in orange) , as predicted by Corollary 3.
@@
In a second simulation study we use system parameter $\theta=0.2$ and set the absorber at $\theta_{\rm abs}=\theta-\delta_{n}$, with displacement $\delta_{n}=n^{-1/2}\tau_{n}$ for $n=6\times 10^{5}$ and $\tau_{n}=7$. We perform the same measurement as above and extract the pattern counts $N_{\alpha,n}(\omega)$ for each trajectory $\omega$. We then use the pattern counts estimator (23) to estimate $\theta$, taking $\tilde{\theta}_{n}=\theta$. This amounts to assuming that the first stage of the general estimation procedure outlined in section VII gives a perfect estimator, which is then used in setting the absorber parameter in the second step. While this procedure cannot be used in a practical situation, the study has theoretical value in that it allows us to study the performance of the pattern counts estimator in its own right, rather than in conjunction with the first step estimator. Figure 3 shows that the final estimator $\hat{\theta}_{n}$ has Gaussian distribution with variance closely matching $1/(nf_{\theta})$, thus achieving the QCRB in this idealised setup. This can also be seen by comparing the “effective” Fisher information $F_{\rm eff}:=(n\overline{(\hat{\theta}_{n}-\theta)^{2}})^{-1}$ where the mean square error is estimated from the data, with the QFI rate $f_{\theta}$; the former is equal to $F_{\rm eff}=13.8$ while the latter is $f_{\theta}=13.5$.
@@
In the third simulations study we implement the full estimation procedure described in section VII including the first stage estimator. In the first step we fix $\theta=0.2$, and run the Markov chain (without the absorber) for $\tilde{n}=4\times 10^{5}$. We then perform sequential measurements in the standard basis on the output noise units to obtain a measurement trajectory from which we compute $\tilde{\theta}_{n}$ by equating the empirical mean (the average number of $1$ counts) with the stationary expected value $c_{\tau}:={\rm Tr}(\rho^{\rm ss}_{\tau}K_{\tau,1}^{*}K_{\tau,1})$, where $K_{\tau,i}$ are the system’s Kraus operators. The effective Fisher information of this estimator is $F_{\rm eff}=4.06$, significantly lower than the QFI $f_{\theta}=13.5$.
@@
From these average counts we compute the estimator (23), where the Fisher information $f$ is computed at $\tilde{\theta}_{n}$ and multipled by $n^{\prime}/n$ to account for the smaller number of samples used in the last step. The results are illustrated in Figure 4 which compares the histogram of $\hat{\theta}_{n}$ (in blue) with the density of a normal distribution with mean $\theta$ and variance $1/(nf_{\theta})$. We find that there is a good fit with the normal distribution but less accurate that that of the second simulation study, cf. Figure 3. This is expected, since the final estimator is based on a two stage estimation process and does not use any prior information about $\theta$. A more accurate measure of the protocol’s performance is given by the effective Fisher information which works out as $F_{\rm eff}=10.8$ compared to the QFI rate $f_{\theta}=13.5$, while the effective Fisher information of the first stage was only $4.06$. These simulation results are in agreement with the theoretical arguments put forward in section VII which indicate that the two stage estimator attains the QCRB asymptotically.
@@ IX Conclusions and Outlook
The estimation strategy consisted of two estimation stages. In the first stage a rough estimator is computed from outcomes of simple output measurements by using a fast but non-optimal procedure, e.g. equating the empirical counts with the expected value at the estimated parameter. In the second stage, we used a coherent quantum absorber Stannigel et al. (2012) to post-process the output Godley and Guţă (2023); Yang et al. (2023). When tuned to the true value of the system parameter, the absorber “reverts” the evolution such that, in the stationary regime, the post-processed output is identical to the input “vacuum” product state. On the other hand, small mismatches between system and absorber parameters lead to slight rotations away from this product state, which can be detected by simple sequential measurements on the output “noise units”. To achieve the perfect “null” setup it would seem natural to use the first stage estimator as absorber parameter, but as shown in Girotti et al. (2024), this leads to non-identifiability issues and sub-optimal final estimators. Instead, we applied the displaced-null technique Girotti et al. (2024) which prescribes an extra parameter shift, calibrated to remove the non-identifiability issue while preserving the optimality of the sequential measurement.
@@
The key theoretical contributions of this work are related to the understanding of the output state and the stochastic measurement process. We introduced the concept of translationally invariant modes (TIMs) of the output and showed how they generate a Bosonic algebra in the asymptotic limit. Each mode is labelled by a binary sequence called a “pattern” and its creation operator is an average of shifted blocks consisting of tensor products operators. We then showed that when the mismatch between system and absorber parameter scales at the estimation rate $n^{-1/2}$, the restriction of the output state to the TIMs becomes a multi-mode coherent state whose displacement depends linearly on the mismatch (quantum Gaussian shift model). Moreover, we showed that the TIMs carry all the quantum Fisher information of the output state and are therefore the relevant quantum statistics of the problem. While homodyne is the standard optimal measurement for such models, in the presence of the additional parameter displacement the modes amplitudes become large and counting measurements become effectively equivalent to homodyne. Surprisingly, we discovered that the sequential counting measurement acts as a joint measurement of all TIMs number operators. Due to the proximity to the vacuum state, typical trajectories consist of a relatively small number of patterns separated by long sequences of 00s. We showed that for large times the patterns counts distribution converges to the Poisson distribution of the TIMs coherent state. These insights allowed us to devise a simple “pattern counting” estimator for estimating the unknown parameter.
@@
Our discrete-time results open the way for fast and optimal continuous-time estimation strategies based on coherent absorber post-processing Yang et al. (2023) and simple pattern counting estimation, as opposed to expensive maximum likelihood methods. Interesting and important topics for future investigations concerns the robustness of the pattern counting method with respect to various types of noises, improving the estimation accuracy for short times, extensions to multi-parameter models and the relationship between the general theory of quantum absorbers Tsang (2024) and quantum estimation.
@@
Acknowledgements. This work was supported by the EPSRC grant EP/T022140/1. MG thanks Juan Garrahan, Mankei Tsang, Dayou Yang, Martin Plenio and Susana Huelga for fruitful discussions. The Python code and the simulation data used in this paper can be found at https://github.com/AGodley/Estimating_QMC_with_CA_and_PatternCounting
@@ Appendix A Displaced-null measurements and asymptotic normality
In this section we review the concept of quantum local asymptotic normality (QLAN) in its simplest form involving a one-parameter family of pure states, and use it to better understand the asymptotic theory of SLD and displaced-null measurements. For the general QLAN theory we refer to Guţă and Kahn (2006); Guţă and Jencova (2007); Kahn and Guţă (2009); Gill and Guţă (2013); Guţă et al. (2008); Butucea et al. (2016) and Yamagata et al. (2013); Fujiwara and Yamagata (2020, 2023). For our purposes, QLAN expresses the fact that the statistical model $|\psi_{\theta}\rangle^{\otimes n}$ describing an ensemble of $n$ identically prepared systems can be approximated by a simpler quantum Gaussian model consisting of a coherent state of a continuous variables (CV) system, whose mean is related to $\theta$ by a linear transformation. Below, we will show that the SLD measurement and the null measurement in the multi-copy model correspond to measuring canonical variables and respectively displaced number operators in the Gaussian picture.
@@
We now construct a Gaussian model which approximates the multiple copies model in the neighbourhood of $\theta_{0}$. Let $\mathcal{F}_{d-1}=\mathcal{F}^{\otimes{(d-1)}}$ be the Fock space of $d-1$ CV modes with annihilation operators $(a_{1},\dots,a_{d-1})$, and let $|{\bf n}\rangle=\otimes_{j=1}^{d-1}|n_{j}\rangle$ be the Fock basis, where ${\bf n}=(n_{1},\dots,n_{d-1})\in\mathbb{N}^{d-1}$. We further denote by $|{\bf z}\rangle=\otimes_{j=1}^{d-1}|z_{i}\rangle$ a coherent state with amplitude ${\bf z}=(z_{1},\dots,z_{d-1})\in\mathbb{C}^{d-1}$ such that $\langle{\bf z}|a_{j}|{\bf z}\rangle=z_{j}$. On the Fock space we consider the quantum statistical model with parameter $u\in\mathbb{R}$ consisting of coherent states
@@
$Q^{n}_{j}$ ::=:= $\frac{1}{\sqrt{2n}}\sum_{i=1}^{n}{(|0\rangle\langle j|+|j\rangle\langle 0|)_{j}}$
$P^{n}_{j}$ ::=:= $\frac{1}{\sqrt{2n}}\sum_{i=1}^{n}{(-i|0\rangle\langle j|+i|j\rangle\langle 0|)_{j}}$ (28)
@@
where $u,v$ are fixed local parameters. The reader will recognise that for $d=2$ these statements are statistical reformulations of results from coherent spin states Radcliffe (1971).
The third type of QLAN convergence is akin to the Le Cam strong convergence of statistical models in classical statistics van der Vaart (1998); it states that the multi-copy and Gaussian models can be mapped into each other operationally by means of quantum channels. Let $\mathcal{S}_{n}$ be the symmetric subspace of $\left(\mathbb{C}^{d}\right)^{\otimes n}$ and let $V_{n}:\mathcal{S}_{n}\to\mathcal{F}_{d-1}$ be the isometry
@@
where $|{\bf n};n\rangle$ is the normalised projection onto ${\cal S}_{n}$ of the basis vector $\left(\otimes_{j=1}^{d-1}|j\rangle^{\otimes n_{j}}\right)\otimes|0\rangle^{n-|{\bf n}|}\in\left(\mathbb{C}^{d}\right)^{\otimes n}$. Then the following holds for any $10<\gamma<$ 1 Butucea et al. (2016)
@@
We now move away from SLD measurements and consider the QLAN perspective on displaced-null measurements. An important point here is that the parameter $u$ of the Gaussian shift model $|{\bf c}u\rangle$ is not completely arbitrary but can be considered to be bounded by $n^{\epsilon}$, where $n$ is the sample size of the qudit ensemble. Indeed, recall that the reference parameter $\theta_{0}$ should be thought of as a preliminary estimator $\tilde{\theta}_{n}$ obtained by measuring a sub-sample of $\tilde{n}=n^{1-\epsilon}$ qudits. Assuming this has standard concentration properties i.e. $|\theta-\tilde{\theta}_{n}|<n^{-1/2+\epsilon}$ with high probability, and writing $\theta=\tilde{\theta}_{n}+u/\sqrt{n}$ we obtain $|u|\leq n^{\epsilon}$ Girotti et al. (2024).

### v3

@@ 7 Pattern counting estimator
    set the absorber parameter by using the displaced-null measurement technique developed in [1] and run the system-absorber dynamics for the remainder of the time;
@@
The first step of the adaptive protocol is to use $\tilde{n}:=n^{1-\epsilon}\ll n$ output units to produce a rough preliminary estimator $\tilde{\theta}_{n}$. This can be done by performing a repeated standard basis measurement on the output (without using an absorber). Typically, the estimator $\tilde{\theta}_{n}$ will have variance scaling with the standard rate $\tilde{n}^{-1}=n^{-1+2\epsilon}$, and $\tilde{n}^{1/2}(\tilde{\theta}_{n}-\theta)$ will satisfy the central limit theorem and a concentration bound ensuring that $|\tilde{\theta}_{n}-\theta|=O(n^{-1/2+\epsilon})$ with high probability. For instance, one can define $\tilde{\theta}_{n}$ by equating the empirical counting rate with the theoretical rate $n_{\tau}:={\rm Tr}(\rho^{\rm ss}_{\tau}K_{\tau,1}^{*}K_{\tau,1})$, see [112, 108].
In the second step we set the absorber at a parameter value $\theta_{0}$ and run the system-absorber quantum Markov chain for the reminder of the time $n^{\prime}=n-\tilde{n}$. The naive choice for $\theta_{0}$ is our best guess $\tilde{\theta}_{n}$ about $\theta$, based on the first stage measurement. However, with this choice, the counting measurement suffers from the non-identifiability issue described in section 2. This can be resolved by further displacing the absorber parameter by an amount $\delta_{n}:=\tau_{n}/\sqrt{n}$ where $\tau_{n}=n^{3\epsilon}$ so that $\theta_{\rm abs}:=\tilde{\theta}_{n}-\delta_{n}$. As usual we write $\theta=\tilde{\theta}_{n}+u_{n}/\sqrt{n}$ where $u_{n}$ is a local parameter satisfying $|u_{n}|\leq n^{\epsilon}$, so that $\theta=\theta_{\rm abs}+(u_{n}+\tau_{n})/\sqrt{n}$.
@@
We now describe the construction of the estimator from the outcomes of this last stage measurement. The estimator will be built using the “pattern extraction” function $\{N_{\alpha,n}(\omega):\alpha\in\mathcal{P}\}\in\mathbb{N}^{\mathcal{P}}$ that we defined in the previous section; we recall that it associates to each trajectory $\omega\in\{0,1\}^{n}$ a set of pattern counts, which is uniquely determined by the condition that $\omega$ is a maximal union of contiguous patterns separated by sequences of 0s of length at least $n^{\gamma}$ with a fixed $0<\gamma<1$. This means that the algorithm will not detect any pattern which contains a sequence of zeros of length larger than $n^{\gamma}$, since this would be seen as being made up of several identified patterns.
For illustration, if we pick $\gamma\leq 0.41$, the sequence in Figure 1 d) would result in the following pattern counting: $3$ patterns of the form $(1)$, 22 patterns of the form $(11)$ and one pattern of the types $(101)$ and $(111)$ (the patterns are colored in red in the picture). Indeed, the sequence has $72$ bits, therefore we would perform a ‘cut’ after 6 consecutive 0s since $72^{\gamma}\approx 5.7$.
Since $\theta-\theta_{\rm abs}$ is of order $n^{-1/2+3\epsilon}$, the expected number of patterns is of order $n^{6\epsilon}$ and the gaps between them are expected to be of size $n^{1-6\epsilon}$. Therefore, we choose the buffer parameter such that $0<\gamma<1-6\epsilon$, which is always possible as long $\epsilon<1/6$.
We now introduce the final estimator using an intuitive argument based on extrapolating the results of Theorem 3 from fixed to slowly growing local parameters $u_{n}+\tau_{n}$. This means that $N_{\alpha,n}(\omega)$ is approximately distributed as ${\rm Poisson}((u_{n}+\tau_{n})^{2}|\mu_{\alpha}|^{2})$, for large $n$; since $\tau_{n}=n^{3\epsilon}$ is larger than $u_{n}=O(n^{\epsilon})$, the intensity of the Poisson distribution diverges with $n$, and the distribution can be approximated further by the normal $N((u_{n}+\tau_{n})^{2}|\mu_{\alpha}|^{2},(u_{n}+\tau_{n})^{2}|\mu_{\alpha}|^{2})$ with the same mean and variance. Using
@@
At the moment we only have a rigorous proof of this statement assuming a stronger version of Proposition 2, which we were not able to obtain; however, we point out that Theorem 2 in [1] establishes a similar optimality result in the case of multi-parameter estimation with independent, identical copies.
@@
The proof of the following theorem and the relationship between the extra hypothesis in Eq. (27) and Proposition 2 can be found in Appendix G. In few words, the difference between the statement of Proposition 2 and the hypothesis in Eq. (27) is the following: Eq. (23) concerns the asymptotic behaviour of the probability of observing a finite number of patterns, while, in order to prove Theorem 4, one would need the asymptotic behaviour of the probability of observing a collections of patterns with a growing number of patters.
@@
8 Numerical experiments
In this section we illustrate the estimation protocol through numerical simulations, using a qubit model inspired by the previous work [1]. The simulations are implemented in Python using the QuTiP package [137].
@@ 8 Numerical experiments
In a first simulation study we verify the predictions made in Corollary 3. We run the dynamics for a total of $n=6\times 10^{5}$ time steps, with absorber parameter $\theta_{\rm abs}=\theta_{0}$ (cf. section 3.1) and system parameter $\theta=\theta_{0}+u/\sqrt{n}$ such that $\theta=0.2$ and the local parameter is $u=2$.
For each run we perform repeated measurements in the standard basis of the output units, to produce a measurement record $\omega=(\omega_{1},\dots,\omega_{n})$. For each such trajectory, the pattern counts $N_{\alpha,n}(\omega)$ are obtained by identifying patterns (sequences starting and ending with a 1) which are separated by at least $n^{\gamma}$ 0s and no pattern contains more than $n^{\gamma}$ 0s, where $\gamma>0$ is a small parameter. This can be done by combing through the sequence and identifying occurrences of a given such pattern padded by $n^{\gamma}$ 0s to the left and right (taking care of the special case of the first and last patterns). Note that for any given $n$ this procedure will not count patterns with more than $n^{\gamma}$ successive 0s. Since the mean counts for each pattern $\alpha$ is $|\mu_{\alpha}|^{2}u^{2}$ and $|\mu_{\alpha}|^{2}$ decays exponentially with $|\alpha|$, we find that patterns of such length are unlikely to occur for large $n$. The results of $N=2000$ independent repetitions of the experiment are illustrated in Figure 2 which shows a good match between the counts histograms corresponding to several patterns (in blue) and the theoretical Poisson distributions ${\rm Poisson}(|\mu_{\alpha}|^{2}u^{2})$ (in orange) , as predicted by Corollary 3.
@@
In a second simulation study we use system parameter $\theta=0.2$ and set the absorber at $\theta_{\rm abs}=\theta-\delta_{n}$, with displacement $\delta_{n}=n^{-1/2}\tau_{n}$ for $n=6\times 10^{5}$ and $\tau_{n}=7$. We perform the same measurement as above and extract the pattern counts $N_{\alpha,n}(\omega)$ for each trajectory $\omega$. We then use the pattern counts estimator (25) to estimate $\theta$, taking $\tilde{\theta}_{n}=\theta$. This amounts to assuming that the first stage of the general estimation procedure outlined in section 7 gives a perfect estimator, which is then used in setting the absorber parameter in the second step. While this procedure cannot be used in a practical situation, the study has theoretical value in that it allows us to study the performance of the pattern counts estimator in its own right, rather than in conjunction with the first step estimator. Figure 3 shows that the final estimator $\hat{\theta}_{n}$ has Gaussian distribution with variance closely matching $1/(nf_{\theta})$, thus achieving the QCRB in this idealised setup. This can also be seen by comparing the “effective” Fisher information $F_{\rm eff}:=(n\overline{(\hat{\theta}_{n}-\theta)^{2}})^{-1}$ where the mean square error is estimated from the data, with the QFI rate $f_{\theta}$; the former is equal to $F_{\rm eff}=13.8$ while the latter is $f_{\theta}=13.5$.
@@
In the third simulations study we implement the full estimation procedure described in section 7 including the first stage estimator. In the first step we fix $\theta=0.2$, and run the Markov chain (without the absorber) for $\tilde{n}=4\times 10^{5}$. We then perform sequential measurements in the standard basis on the output noise units to obtain a measurement trajectory from which we compute $\tilde{\theta}_{n}$ by equating the empirical mean (the average number of $1$ counts) with the stationary expected value $c_{\tau}:={\rm Tr}(\rho^{\rm ss}_{\tau}K_{\tau,1}^{*}K_{\tau,1})$, where $K_{\tau,i}$ are the system’s Kraus operators. The effective Fisher information of this estimator is $F_{\rm eff}=4.06$, significantly lower than the QFI $f_{\theta}=13.5$.
@@
From these average counts we compute the estimator (25), where the Fisher information $f$ is computed at $\tilde{\theta}_{n}$ and multipled by $n^{\prime}/n$ to account for the smaller number of samples used in the last step. The results are illustrated in Figure 4 which compares the histogram of $\hat{\theta}_{n}$ (in blue) with the density of a normal distribution with mean $\theta$ and variance $1/(nf_{\theta})$. We find that there is a good fit with the normal distribution but less accurate that that of the second simulation study, cf. Figure 3. This is expected, since the final estimator is based on a two stage estimation process and does not use any prior information about $\theta$. A more accurate measure of the protocol’s performance is given by the effective Fisher information which works out as $F_{\rm eff}=10.8$ compared to the QFI rate $f_{\theta}=13.5$, while the effective Fisher information of the first stage was only $4.06$. These simulation results are in agreement with the theoretical arguments put forward in section 7 which indicate that the two stage estimator attains the QCRB asymptotically.
@@ 9 Conclusions and Outlook
The estimation strategy consisted of two estimation stages. In the first stage a rough estimator is computed from outcomes of simple output measurements by using a fast but non-optimal procedure, e.g. equating the empirical counts with the expected value at the estimated parameter. In the second stage, we used a coherent quantum absorber [116] to post-process the output [115, 2]. When tuned to the true value of the system parameter, the absorber “reverts” the evolution such that, in the stationary regime, the post-processed output is identical to the input “vacuum” product state. On the other hand, small mismatches between system and absorber parameters lead to slight rotations away from this product state, which can be detected by simple sequential measurements on the output “noise units”. To achieve the perfect “null” setup it would seem natural to use the first stage estimator as absorber parameter, but as shown in [1], this leads to non-identifiability issues and sub-optimal final estimators. Instead, we applied the displaced-null technique [1] which prescribes an extra parameter shift, calibrated to remove the non-identifiability issue while preserving the optimality of the sequential measurement.
@@
The key theoretical contributions of this work are related to the understanding of the output state and the stochastic measurement process. We introduced the concept of translationally invariant modes (TIMs) of the output and showed how they generate a Bosonic algebra in the asymptotic limit. Each mode is labelled by a binary sequence called a “pattern” and its creation operator is an average of shifted blocks consisting of tensor products operators. We then showed that when the mismatch between system and absorber parameter scales at the estimation rate $n^{-1/2}$, the restriction of the output state to the TIMs becomes a multi-mode coherent state whose displacement depends linearly on the mismatch (quantum Gaussian shift model). Moreover, we showed that the TIMs carry all the quantum Fisher information of the output state and are therefore the relevant quantum statistics of the problem. While homodyne is the standard optimal measurement for such models, in the presence of the additional parameter displacement the modes amplitudes become large and counting measurements become effectively equivalent to homodyne. Surprisingly, we discovered that the sequential counting measurement acts as a joint measurement of all TIMs number operators. Due to the proximity to the vacuum state, typical trajectories consist of a relatively small number of patterns separated by long sequences of 0s. We showed that for large times the patterns counts distribution converges to the Poisson distribution of the TIMs coherent state. These insights allowed us to devise a simple “pattern counting” estimator for estimating the unknown parameter.
@@
Our discrete-time results open the way for fast and optimal continuous-time estimation strategies based on coherent absorber post-processing [2] and simple pattern counting estimation, as opposed to expensive maximum likelihood methods. Interesting and important topics for future investigations concerns the robustness of the pattern counting method with respect to various types of noises, improving the estimation accuracy for short times, extensions to multi-parameter models and the relationship between the general theory of quantum absorbers [136] and quantum estimation.
@@
Acknowledgements. This work was supported by the EPSRC grant EP/T022140/1. MG thanks Juan Garrahan, Mankei Tsang, Dayou Yang, Martin Plenio and Susana Huelga for fruitful discussions. The Python code and the simulation data used in this paper can be found at https://www.doi.org/10.5281/zenodo.16755954
@@ Appendix A Displaced-null measurements and asymptotic normality
In this section we review the concept of quantum local asymptotic normality (QLAN) in its simplest form involving a one-parameter family of pure states, and use it to better understand the asymptotic theory of SLD and displaced-null measurements. For the general QLAN theory we refer to [117, 121, 118, 122, 119, 120] and [123, 124, 125]. For our purposes, QLAN expresses the fact that the statistical model $|\psi_{\theta}\rangle^{\otimes n}$ describing an ensemble of $n$ identically prepared systems can be approximated by a simpler quantum Gaussian model consisting of a coherent state of a continuous variables (CV) system, whose mean is related to $\theta$ by a linear transformation. Below, we will show that the SLD measurement and the null measurement in the multi-copy model correspond to measuring canonical variables and respectively displaced number operators in the Gaussian picture.
@@
We now construct a Gaussian model which approximates the multiple copies model in the neighbourhood of $\theta_{0}$. Let $\mathcal{F}_{d-1}=\mathcal{F}^{\otimes{(d-1)}}$ be the Fock space of d1d-1 CV modes with annihilation operators $(a_{1},\dots,a_{d-1})$, and let $|{\bf n}\rangle=\otimes_{j=1}^{d-1}|n_{j}\rangle$ be the Fock basis, where ${\bf n}=(n_{1},\dots,n_{d-1})\in\mathbb{N}^{d-1}$. We further denote by $|{\bf z}\rangle=\otimes_{j=1}^{d-1}|z_{i}\rangle$ a coherent state with amplitude ${\bf z}=(z_{1},\dots,z_{d-1})\in\mathbb{C}^{d-1}$ such that $\langle{\bf z}|a_{j}|{\bf z}\rangle=z_{j}$. On the Fock space we consider the quantum statistical model with parameter $u\in\mathbb{R}$ consisting of coherent states
@@
$Q^{n}_{j}$ $:=$ $\frac{1}{\sqrt{2n}}\sum_{i=1}^{n}{(|0\rangle\langle j|+|j\rangle\langle 0|)_{j}}$
$P^{n}_{j}$ $:=$ $\frac{1}{\sqrt{2n}}\sum_{i=1}^{n}{(-i|0\rangle\langle j|+i|j\rangle\langle 0|)_{j}}$ (30)
@@
where $u,v$ are fixed local parameters. The reader will recognise that for $d=2$ these statements are statistical reformulations of results from coherent spin states [138].
The third type of QLAN convergence is akin to the Le Cam strong convergence of statistical models in classical statistics [139]; it states that the multi-copy and Gaussian models can be mapped into each other operationally by means of quantum channels. Let $\mathcal{S}_{n}$ be the symmetric subspace of $\left(\mathbb{C}^{d}\right)^{\otimes n}$ and let $V_{n}:\mathcal{S}_{n}\to\mathcal{F}_{d-1}$ be the isometry
@@
where $|{\bf n};n\rangle$ is the normalised projection onto ${\cal S}_{n}$ of the basis vector $\left(\otimes_{j=1}^{d-1}|j\rangle^{\otimes n_{j}}\right)\otimes|0\rangle^{n-|{\bf n}|}\in\left(\mathbb{C}^{d}\right)^{\otimes n}$. Then the following holds for any $0<\gamma<1$ [120]
@@
We now move away from SLD measurements and consider the QLAN perspective on displaced-null measurements. An important point here is that the parameter $u$ of the Gaussian shift model $|{\bf c}u\rangle$ is not completely arbitrary but can be considered to be bounded by $n^{\epsilon}$, where $n$ is the sample size of the qudit ensemble. Indeed, recall that the reference parameter $\theta_{0}$ should be thought of as a preliminary estimator $\tilde{\theta}_{n}$ obtained by measuring a sub-sample of $\tilde{n}=n^{1-\epsilon}$ qudits. Assuming this has standard concentration properties i.e. $|\theta-\tilde{\theta}_{n}|<n^{-1/2+\epsilon}$ with high probability, and writing $\theta=\tilde{\theta}_{n}+u/\sqrt{n}$ we obtain $|u|\leq n^{\epsilon}$ [1].